x(x-2)^3-35(x-2)^2=0
Answer by mananth(16946)

x(x-2)^3-35(x-2)^2=0
(x-2)^2{x(x-2) -35)}=0
(x-2)^2 { x^2-2x-35}=0
(x-2)^2{x^2—7x+5x-35}=0
(x-2)^2{x(x-7)+5(x-7)}= 0
(x-2)^2{(x-7)(x+5)}=0
(x-2)^2(x-7)(x+5)=0